OpenStudy (anonymous):

Part 1: Explain, in complete sentences, whether or not the expression x2 – 121 fits one of the special patterns. If so, which one? (2 points) Part 2: Explain how the polynomial is factored. (2 points) Part 3: Provide the factors. (2 points)

OpenStudy (anonymous):

x2 - 121 fits the special pattern a^2 - b^2 = (a-b)(a+b), with a = x and b = 11. Therefore, the factorization is (x - 11)(x + 11)